# Reconfigurable Intelligent Surface Aided Vehicular Edge Computing: Joint Phase-shift Optimization and Multi-User Power Allocation

### v1

@@ V-C Testing Stage
The testing phase omits the updating of various network parameters and gradients, and the optimal phase-shift coefficients and power allocation scheme will be obtained from the trained model. The pseudocode of the testing stage is shown in Algorithm 3.
@@ V-D Complexity Analysis
    DDPG: $O\left({(E\cdot S-I)({G_{A}}+{G_{C}}+2{U_{A}}+2{U_{C}})}\right)$
@@
    TD3: $O\left({(E\cdot S-I)({G_{A}}+{G_{C}}+3{U_{A}}+3{U_{C}})}\right)$
@@
    $\begin{gathered}\overbrace{O\left({K\cdot(E\cdot S-I)({G_{A}}+{G_{C}}+2{U_{A}}+2{U_{C}})}\right)}^{Local{\text{ }}agent}+\hfill\\ \underbrace{O\left({(E\cdot S-I)(2{G_{C}}+4{U_{C}})}\right)}_{Global{\text{ }}critic}\hfill\\ \end{gathered}$
@@ VI Simulation Results
In this section, we perform simulation experiments to validate our proposed SARL-MARL scheme. The simulation tool is Python 3.9. To verify the performance of the proposed scheme, we compare it with the traditional centralized DDPG and TD3 algorithm. It is assumed that the vehicle is travelling on an urban road and the position coordinates of the BS are $(0,0,25)$ and the position coordinates of the RIS are (250, 220, 25). $8$ vehicles are involved in the experiment, each with a randomly chosen speed within $10$ to $15$ m/s. The learning parameters of the DDPG scheme in Algorithm 1 and the modified MADDPG scheme in Algorithm 2 are basically same. The numbers of hidden layers for actor and critic networks are 2 and 3, respectively. The size of hidden layers for actor network is 512 and 256, and the size of hidden layers for critic network is 1024, 512 and 256. The key parameters are shown in Table I.
@@
Fig. 4 presents the comparison between the proposed scheme and the TD3 and DDPG schemes in terms of reward convergence performance. The proposed scheme exhibits very stable convergence performance with less oscillations compared to the fully centralized TD3 and DDPG.TD3 and DDPG, as fully centralized approaches, require the observations and actions of all agents as inputs, which not only limits their cap ability to efficiently solve the individual performance problem, which may lead to sub-optimal rewards and unstable convergence situations, but also puts a base station server imposes a significant information burden. On the contrary, the proposed framework combines the advantages of the SARL and MARL algorithms and exhibits superior convergence performance. It is capable of effectively selecting local and global reward maximization strategies for all agents, thus promoting cooperation among agents and further improving the overall system performance.
@@
Fig. 6 illustrates the variation of power consumption of each VU at different task arrival rates. It can be observed that the power consumption of our proposed SARL-MARL scheme increases with the task arrival rate. In contrast, the schemes without RIS assistance and with random RIS phase-shift have their power consumption slightly higher, highlighting the advantage of the DDPG algorithm for training the RIS phase-shifting coefficients. Finally, due to the relative advantage of the improved MADDPG algorithm, all these schemes consume significantly less power than the DDPG and TD3 schemes. Since DDPG and TD3 are fully centralized methods, they may fall into sub-optimal solutions and thus fail to achieve a good power allocation scheme.
Fig. 7 shows the variation of buffer lengths for each VU at different task arrival rates. As the task arrival rate increases, the buffer lengths of all VUs also increase. In particular, the buffer length of the DDPG algorithm is relatively low because it consumes too much power to transmit buffer tasks. On the contrary, the TD3 algorithm fails to balance power consumption and buffer length effectively, leading to task stacking at larger task arrival rates. At time slot $t$, when the task arrival rate is 4M/bps, the instantaneous buffer length of the TD3 algorithm is truncated to 50Kb, but the actual buffer length pileup reaches 750 Kb. Our proposed scheme exhibits a relatively small buffer length with minimal power consumption, successfully balancing the relationship between power and buffer length.

### v2

@@ V-C Testing Stage
The testing phase omits the updating of various network parameters and gradients, and the optimal phase-shift coefficients and power allocation scheme will be obtained from the trained model. We conduct tests based on the trained model, where each round contains 50 test slots, and each result is obtained by averaging the results of sequential 10 rounds. The pseudocode of the testing stage is shown in Algorithm 3.
@@ V-D Complexity Analysis
    DDPG: $\mathcal{O}\left({(E\cdot S-I)({G_{A}}+{G_{C}}+2{U_{A}}+2{U_{C}})}\right)$
@@
    TD3: $\mathcal{O}\left({(E\cdot S-I)({G_{A}}+{G_{C}}+3{U_{A}}+3{U_{C}})}\right)$
@@
    $\begin{gathered}\overbrace{\mathcal{O}\left({K\cdot(E\cdot S-I)({G_{A}}+{G_{C}}+2{U_{A}}+2{U_{C}})}\right)}^{Local{\text{ }}agent}+\hfill\\ \underbrace{\mathcal{O}\left({(E\cdot S-I)(2{G_{C}}+4{U_{C}})}\right)}_{Global{\text{ }}critic}\hfill\\ \end{gathered}$
@@ VI Simulation Results
In this section, we perform simulation experiments to validate our proposed scheme. The simulation tool is Python 3.9. To verify the performance of the proposed scheme, we compare it with the traditional centralized DDPG and TD3 algorithm. It is assumed that the vehicle is travelling on an urban road and the position coordinates of the BS are $(0,0,25)$ and the position coordinates of the RIS are (250, 220, 25). $8$ vehicles are involved in the experiment, each with a randomly chosen speed within $10$ to $15$ m/s. The learning parameters of the DDPG scheme in Algorithm 1 and the modified MADDPG scheme in Algorithm 2 are same. The numbers of hidden layers for actor and critic networks are 2 and 3, respectively. The size of hidden layers for actor network is 512 and 256, and the size of hidden layers for critic network is 1024, 512 and 256. The key parameters are shown in Table I.
@@
Fig. 4 presents the comparison between the proposed scheme and the TD3 and DDPG schemes in terms of reward convergence performance. The proposed scheme exhibits very stable convergence performance with less oscillations compared to the fully centralized TD3 and DDPG. TD3 and DDPG, as fully centralized approaches, require the observations and actions of all agents as inputs, which not only limits their cap ability to efficiently solve the individual performance problem. In such a way, it may lead to sub-optimal rewards and unstable convergence situations, and also puts a BS server imposes a significant information burden. On the contrary, the proposed framework combines the advantages of the SARL and MARL algorithms and exhibits superior convergence performance. It is capable of effectively selecting local and global reward maximization strategies for all agents, thus promoting cooperation among agents and further improving the overall system performance.
@@
[…]